1. IDL PROCEDURES—SPECIAL FOR FRANK!

1.1. For Controlling the Dish

1. pointdish moves the telescope to a specified (alt enc,az enc). You can specify either one or both coordinates. Example:

print, pointdish( alt=20)

pointdish has the /nolimits option, which ignores the software limits. Useful for testing, but dangerous if you make a mistake; you can ruin the optical fiber and the drive system.

2. pointl moves the telescope to a specified (alt,az). You must specify both coordinates. Example:

print, pointl( alt=20, az=180)

It points to earth-based coordinates unless the /encoder keyword is set.

3. maint moves the telescope to maintenance position

4. stow moves the telescope to stow position

5. pointl radec moves the telescope to a specified (ra,dec), sun, or moon, depending on which keywords you set.

6. follow tracks a given (ra,dec), sun, or moon and can write a file containing relevant pointing data.

1.2. For Taking Data

1. noise turns on or off the calibration noise diode. Examples: noise, /on and noise, /off

2. rxpwr returns the channel-integrated spectral power from the ROACH box as input. So to see the rf power of the noise diode:

noise, /off
wait,4
print, rxpwr()
noise, /on
wait,4
print, rxpwr()

3. set lhp sets the frequency and level of the l.o. Important: the resolution of the l.o. is 10 KHz; see the documentation for this procedure.

4. startchart1 runs a ‘chart recorder’ using the channel-integrated spectral power from the ROACH box as input.

5. getspect sets the l.o. and the noise diode, then averages a specified number of spectra from the ROACH box and optionally plots or saves the result in an IDL save file, which contain ancillary info such as Julian day. ra, dec, etc.

6. leuschner rx Take and record spectra with the ROACH box.

7. filename generate a filename of the form: jjjjjjj.jjjjjj sn0000 nd1 1270123456

8. fits to sav converts a fits file, which contains nrspectra spectra, into a sav file with a single spectrum equal to the avg, also the median. Also lots of ancillary info, all in a structure

9. frq calculate the array of IF or RF freqs from the tags in the standard output structure.